A Chemist has 35% NaCl solution and 50% NaCl solution. How many liters of each should he mix to get 10L of NaCl solution with 40% base concentration?
Please help with this word problem involving systems of linear equations.

Answer by fractalier(6550)

Call the amount of 35% solution, x.
Then the amount of 50% solution would be 10-x.
The set up looks like this:
.35x + .50(10-x) = .40(10)
Now work it out...
.35x + 5 - .50x = 4
-.15x = -1
x = 6.67 L of 35% and then
10-x = 3.33 L of 50% solution